The product of 8 in the sum of 4 and a number is 112. What is the number?
The quotient of a number and 3 is the same as the difference of the number doubled and 5. What is the number?
The sum of three numbers is 69. If the second number is equal to the first diminished by 8, and the third number is 5 times the first, what are the numbers?





Answer by stanbon(48550)

The product of 8 in the sum of 4 and a number is 112. What is the number?
8(4+x)=112
32+8x=112
8x=80
x=10
----------------
The quotient of a number and 3 is the same as the difference of the number doubled and 5. What is the number?
x/3 = 2x-5
Multiply thru by 3 to get:
x=2x^2-5x
2x^2-6x=0
2x(x-3)=0
x=0 or x=3
--------------
The sum of three numbers is 69. If the second number is equal to the first diminished by 8, and the third number is 5 times the first, what are the numbers?
Let the 1st be "x".
The second is x-8
The third is 5x
EQUATION:
x+x-8+5x=69
7x-8=69
7x=77
x=11
----------
1st=x=11
2nd=x-8=11-8=3
3rd=5x=55
Cheers,
Stan H.